OpenStudy (anonymous):

Find an equation of the line that passes through the given points. (Let x be the independent variable and y be the dependent variable.) (2, 8) and (3, 15)

OpenStudy (anonymous):

over 1 up 7 slope is 7

OpenStudy (anonymous):

then use \[y-y_1=m(x-x_1)\] with ,\[x_1=2.y_1=8,m=7\]to find your equation

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

-8=7x-14